OpenStudy (anonymous):

Seperate 116 into two parts, one of which shall be 3 times the other. Show ALL steps

hero (hero):

x + y = 116 x = 3y

hero (hero):

I was already helping her estudier...

OpenStudy (anonymous):

x +3x =116 -> x= 116/4 = 29 and 87

OpenStudy (anonymous):

I like getting opinions or answers from more than one person to ensure rightness

OpenStudy (anonymous):

@Hero Sorry, I saw the post and didn't understand...

hero (hero):

You saw my post and didn't understand....You just did the same exact steps

hero (hero):

If you substitute 3y for x you get 3y + y = 4y

OpenStudy (anonymous):

i like to do these as percentages. so x= 75% of 116, and y =25% of 116. Not much difference, but i like it :-)

hero (hero):

set 4y = 116 and solve for x

hero (hero):

I mean for y

OpenStudy (anonymous):

@ Hero My steps are not the same and your post was incomprehensible. Put your answer in one post in future. Also don't be an retrice I gave u a medal for your lousy answer.

hero (hero):

It wasn't a lousy answer. Unlike you, I give helpful hints to try to lead the person to the answer...Not just post the answer all in one gulp...And by the way, I'm not an as*

hero (hero):

I translated the word problem into algebra. That's all I did

OpenStudy (anonymous):

Stop fighting like children (unless you are children of course)

hero (hero):

You must have never heard of systems of equations before

hero (hero):

I didn't know that systems of equations were "incomprehensible"

OpenStudy (anonymous):

The answer fits on one line, if u truly wish to guide the student as u put it, then I suggest u check back in each post whether or not the student understood up to that point. I apologized for not understanding that u were making another post (I am not psychic). And I gave u a medal for your incomplete answer. I don't then expect to be berated by a fool for my gentility.

hero (hero):

You just called me a fool, but you want me to be convinced that you're not one? Once again, I don't think it was difficult to understand that I was trying to give the student ideas about how to approach solving the problem. For some reason, instead of me explaining my approach here, you think I was trying to be rude to you. I didn't call you a fool or say that what you did was 'incomprehensible'. Your misunderstanding of my approach however is 'incomprehensible'. I personally know that the moderators prefer if you help lead the student to the answer rather than just post them. I don't know why you're being mean.

OpenStudy (anonymous):

I'm not trying to convince u of anything, if u have a problem with anything I did, take it up with the moderators.

hero (hero):

Nah, I don't have to do anything. The situation will take care of itself..